# Solution of some tiling open problems of Propp, Lai, and some related results

### v2

@@ 7. Aztec rectangles with collinear unit holes
Unless $m=n$ (when it becomes the Aztec diamond $AD_{n}$), the Aztec rectangle $AR(m,n)$ does not have any domino tilings: it is not hard to check that if we color the square grid in chessboard fashion, $AR(m,n)$ has $n-m$ more unit squares of one color than the other. This can be fixed as follows. Label the bottom-most unit squares in $AR(m,n)$ by $1,\ldots,n$ from left to right. For any subset $S\subseteq[n]\coloneqq\{1,\ldots,n\}$ of size $m$, we delete the unit squares labeled with elements of $[n]\setminus S$ from $AR(m,n)$, and denote the resulting region by $AR(m,n;S)$ (see the picture on the left in Figure 34 for an illustration). Then the region $AR(m,n;S)$ always admits domino tilings; the number of them is given by (see [4][16][28])
@@
Next, for any subset $T=\{t_{1},\ldots,t_{m}\}\subseteq\{n+1\}$ of size $m$ (again, with elements of $T$ written in increasing order), consider the region $\overline{AR}(m,n;T)$ defined as follows. From $AR(m,n)$, remove the $n$ unit squares at the bottom. Note that the resulting region has $n+1$ unit squares on the (new) bottom. Label these unit squares by $11,\ldots,n+$ 1, from left to right. Finally, delete the unit squares labeled by elements of $T$, and denote the resulting region by $\overline{AR}(m,n;T)$. The number of domino tilings for this region was studied in [16] and [19], and is given by
@@
In [4], the second author proved that if an arbitrary collection of unit squares is deleted from the horizontal symmetry axis of an Aztec rectangle, the number of domino tilings of the resulting region is given by a simple product formula. More precisely, let $m$ and $N$ be positive integers with $2m\leq N$. Then the number of domino tilings of a $2m\times N$ Aztec rectangle, where all the unit squares on the central horizontal row, except the $t_{1}$-st, the $t_{2}$-nd,$\ldots$, and the $t_{2m}$-th unit square have been removed equals
@@
We first prove Theorem 7.1. To do that, we add one layer at the bottom of the region by adding $12N+$ 1 unit squares that form a zigzag path, and then we delete the rightmost newly added unit square (see the left picture in Figure 36). Then, the dual graph of the resulting region is a symmetric graph with a vertex removed from its boundary, to which Theorem 2.1(b) can be applied. If we apply Theorem 2.1(b) to this dual graph, take the dual again, and delete the forced dominos, we get that the number of domino tilings of the region in Theorem 7.1 can be expressed as follows:
@@
The same approach works for proving Theorem 7.2. We add one layer at the bottom of the Aztec rectangle by including $12N+$ 1 unit squares in the same way as in the proof of Theorem 7.1, and then remove the rightmost of these (see the picture on the right in Figure 36). We can again apply Theorem 2.1(b), and after removing the forced forced dominos of the resulting regions, we obtain that the number sought in Theorem 7.2 equals
@@
Let $m$ and $N$ be positive integers with $2m+2\leq N$. Then the number of domino tilings of a $(2m+2)\times N$ Aztec rectangle, where all the unit squares on the horizontal row that is by $2$ below the central row, except for the $t_{1}$-st, the $t_{2}$-nd,$\ldots$, and the $t_{2m+2}$-th unit square, have been removed, equals
@@
Let $R_{0}(2m,N;\{t_{1},\ldots,t_{2m}\})$, $R_{1}(2m+1,N;\{t_{1},\ldots,t_{2m+1}\})$, and $R_{2}(2m+2,N;\{t_{1},\ldots,t_{2m+2}\})$ be the regions described just before equation (7.3), in Theorem 7.1, and in Theorem 7.3, respectively. Under this notation, we want to show that the number of domino tilings of $R_{2}(2m+2,N;\{t_{1},\ldots,t_{2m+2}\})$ is given by (7.8). Note that the region $R_{2}(2m+2,N;\{t_{1},\ldots,t_{2m+2}\})$ has $N-(2m+2)$ unit square holes in it. This is important because we will use an induction on this quantity.
@@
One can check that the expressions on the right hand sides of equations (7.3), (7.4) and (7.8) satisfy the above recurrence relation (7.9). Using this, we will give an inductive proof of Theorem 7.3. We start the proof with an induction on the number of holes in the region $R_{2}(2m+2,N;\{t_{1},\ldots,t_{2m+2}\})$, which is $N-(2m+2)$. Let us call this the outer induction.
@@
When $N-(2m+2)=0$, $\{t_{1},\ldots,t_{2m+2}\}=[2m+2]$ (where $[n]:=\{1,2,\dotsc,n\}$) and the region $R_{2}(2m+2,N;\{t_{1},\ldots,t_{2m+2}\})=R_{2}(2m+2,2m+2;[2m+2])$ becomes the Aztec diamond of order $22m+$ 2. One can check that if we set $N=2m+2$ and $\{t_{1},\ldots,t_{2m+2}\}=[2m+2]$ in the expression on the right hand side of (7.8), we get $2^{\frac{(2m+2)(2m+3)}{2}}$, which is the number of domino tilings of the Aztec diamond of order $22m+$ 2. Thus, the theorem is verified for the case when $N-(2m+2)=0$.
@@
Suppose Theorem 7.3 holds whenever $N-(2m+2)<k$ for some positive integer $k$. Under this assumption, we need to show that the theorem still holds when $N-(2m+2)=k$. To show this, we need another induction. This time we induct on the smallest element of $[N]\setminus\{t_{1},\ldots,t_{2m+2}\}$. We call this the inner induction. If the smallest element is $1$, then the unit square labeled $1$ was removed. But this causes several forced dominos on the left side of the region. Due to this, in this case, the region $R_{2}(2m+2,N;\{t_{1},\ldots,t_{2m+2}\})$ can be identified with $R_{2}(2m+2,N-1;\{t_{1}-1,\ldots,t_{2m+2}-1\})$, which is the region whose number of domino tilings is given by (7.8) by the outer induction hypothesis. Thus we have checked the case when the smallest element of $[N]\setminus\{t_{1},\ldots,t_{2m+2}\}$ is $1$.
@@
Now suppose that Theorem 7.3 holds when $N-(2m+2)=k$ and the smallest element of $[N]\setminus\{t_{1},\ldots,t_{2m+2}\}$ is less than or equal to a positive integer $l$. Under this assumption, we need to show that the theorem still holds when $N-(2m+2)=k$ and the smallest element of $[N]\setminus\{t_{1},\ldots,t_{2m+2}\}$ is $l+1$. Consider the recurrence relation (7.9), and focus on the two terms containing an $R_{2}$-region. It is not difficult to check that when $t_{1}=1$ (which is the case, as we can now assume that the smallest element of $[N]\setminus\{t_{1},\ldots,t_{2m+2}\}$ is $>1$), the smallest element of $N\setminus\{t_{1},\ldots,t_{2m+2}\}$ is greater than that of $N\setminus\{t_{2}-1,\ldots,t_{2m+2}-1,N\}$, and it exceeds the latter by exactly one. Since we already checked that (7.9) is satisfied by (7.3), (7.4), and (7.8), we can conclude that Theorem 7.3 also holds when $N-(2m+2)=k$ and the smallest element of $N\setminus\{t_{1},\ldots,t_{2m+2}\}$ is $l+1$. Thus, the inner induction is checked, and it follows that Theorem 7.3 always holds whenever $N-(2m+2)=k$. Therefore the induction step is verified for the outer induction, and it follows that Theorem 7.3 holds whenever $N-(2m+2)$ is a non-negative integer. This completes the proof.
@@
Acknowledgments. Part of this research was performed while the first author was visiting the Institute for Pure and Applied Mathematics (IPAM), which is supported by the National Science Foundation (Grant No. DMS-1925919). He would like to thank the program organizers and IPAM staff members for their hospitality during his stay in Los Angeles.

### v3

@@ 7. Aztec rectangles with collinear unit holes
Unless $m=n$ (when it becomes the Aztec diamond $AD_{n}$), the Aztec rectangle $AR(m,n)$ does not have any domino tilings: it is not hard to check that if we color the square grid in chessboard fashion, $AR(m,n)$ has nmn-m more unit squares of one color than the other. This can be fixed as follows. Label the bottom-most unit squares in $AR(m,n)$ by $1,\ldots,n$ from left to right. For any subset $S\subseteq[n]\coloneqq\{1,\ldots,n\}$ of size $m$, we delete the unit squares labeled with elements of $n]\setminus S$ from $AR(m,n)$, and denote the resulting region by $AR(m,n;S)$ (see the picture on the left in Figure 36 for an illustration). Then the region $AR(m,n;S)$ always admits domino tilings; the number of them is given by (see [4][16][28])
@@
Next, for any subset $T=\{t_{1},\ldots,t_{m}\}\subseteq\{n+1\}$ of size $m$ (again, with elements of $T$ written in increasing order), consider the region $\overline{AR}(m,n;T)$ defined as follows. From $AR(m,n)$, remove the $n$ unit squares at the bottom. Note that the resulting region has $n+1$ unit squares on the (new) bottom. Label these unit squares by $1,\ldots,n+1$, from left to right. Finally, delete the unit squares labeled by elements of $T$, and denote the resulting region by $\overline{AR}(m,n;T)$. The number of domino tilings for this region was studied in [16] and [19], and is given by
@@
[…]
@@
We first prove Theorem 7.1. To do that, we add one layer at the bottom of the region by adding $2N+1$ unit squares that form a zigzag path, and then we delete the rightmost newly added unit square (see the left picture in Figure 38). Then, the dual graph of the resulting region is a symmetric graph with a vertex removed from its boundary, to which Theorem 2.1(b) can be applied. If we apply Theorem 2.1(b) to this dual graph, take the dual again, and delete the forced dominos, we get that the number of domino tilings of the region in Theorem 7.1 can be expressed as follows:
@@
The same approach works for proving Theorem 7.2. We add one layer at the bottom of the Aztec rectangle by including $2N+1$ unit squares in the same way as in the proof of Theorem 7.1, and then remove the rightmost of these (see the picture on the right in Figure 38). We can again apply Theorem 2.1(b), and after removing the forced forced dominos of the resulting regions, we obtain that the number sought in Theorem 7.2 equals
@@
[…]
@@
Let $R_{0}(2m,N;\{t_{1},\ldots,t_{2m}\})$, $R_{1}(2m+1,N;\{t_{1},\ldots,t_{2m+1}\})$, and $R_{2}(2m+2,N;\{t_{1},\ldots,t_{2m+2}\})$ be the regions described just before equation (7.3), in Theorem 7.1, and in Theorem 7.3, respectively. Under this notation, we want to show that the number of domino tilings of $R_{2}(2m+2,N;\{t_{1},\ldots,t_{2m+2}\})$ is given by (7.8). Note that the region $R_{2}(2m+2,N;\{t_{1},\ldots,t_{2m+2}\})$ has N(2m+2)N-(2m+2) unit square holes in it. This is important because we will use an induction on this quantity.
@@
One can check that the expressions on the right hand sides of equations (7.3), (7.4) and (7.8) satisfy the above recurrence relation (7.9). Using this, we will give an inductive proof of Theorem 7.3. We start the proof with an induction on the number of holes in the region $R_{2}(2m+2,N;\{t_{1},\ldots,t_{2m+2}\})$, which is N(2m+2)N-(2m+2). Let us call this the outer induction.
@@
When N(2m+2)=0N-(2m+2)=0, $\{t_{1},\ldots,t_{2m+2}\}=[2m+2]$ (where $n]:=\{1,2,\dotsc,n\}$) and the region $R_{2}(2m+2,N;\{t_{1},\ldots,t_{2m+2}\})=R_{2}(2m+2,2m+2;[2m+2])$ becomes the Aztec diamond of order $2m+2$. One can check that if we set $N=2m+2$ and $\{t_{1},\ldots,t_{2m+2}\}=[2m+2]$ in the expression on the right hand side of (7.8), we get $2^{\frac{(2m+2)(2m+3)}{2}}$, which is the number of domino tilings of the Aztec diamond of order $2m+2$. Thus, the theorem is verified for the case when N(2m+2)=0N-(2m+2)=0.
@@
Suppose Theorem 7.3 holds whenever N(2m+2)<kN-(2m+2)<k for some positive integer $k$. Under this assumption, we need to show that the theorem still holds when N(2m+2)=kN-(2m+2)=k. To show this, we need another induction. This time we induct on the smallest element of $N]\setminus\{t_{1},\ldots,t_{2m+2}\}$. We call this the inner induction. If the smallest element is $1$, then the unit square labeled $1$ was removed. But this causes several forced dominos on the left side of the region. Due to this, in this case, the region $R_{2}(2m+2,N;\{t_{1},\ldots,t_{2m+2}\})$ can be identified with $R_{2}(2m+2,N-1;\{t_{1}-1,\ldots,t_{2m+2}-1\})$, which is the region whose number of domino tilings is given by (7.8) by the outer induction hypothesis. Thus we have checked the case when the smallest element of $N]\setminus\{t_{1},\ldots,t_{2m+2}\}$ is $1$.
@@
Now suppose that Theorem 7.3 holds when N(2m+2)=kN-(2m+2)=k and the smallest element of $N]\setminus\{t_{1},\ldots,t_{2m+2}\}$ is less than or equal to a positive integer $l$. Under this assumption, we need to show that the theorem still holds when N(2m+2)=kN-(2m+2)=k and the smallest element of $N]\setminus\{t_{1},\ldots,t_{2m+2}\}$ is $l+1$. Consider the recurrence relation (7.9), and focus on the two terms containing an $R_{2}$-region. It is not difficult to check that when $t_{1}=1$ (which is the case, as we can now assume that the smallest element of $N]\setminus\{t_{1},\ldots,t_{2m+2}\}$ is >1>1), the smallest element of $N\setminus\{t_{1},\ldots,t_{2m+2}\}$ is greater than that of $N\setminus\{t_{2}-1,\ldots,t_{2m+2}-1,N\}$, and it exceeds the latter by exactly one. Since we already checked that (7.9) is satisfied by (7.3), (7.4), and (7.8), we can conclude that Theorem 7.3 also holds when N(2m+2)=kN-(2m+2)=k and the smallest element of $N\setminus\{t_{1},\ldots,t_{2m+2}\}$ is $l+1$. Thus, the inner induction is checked, and it follows that Theorem 7.3 always holds whenever N(2m+2)=kN-(2m+2)=k. Therefore the induction step is verified for the outer induction, and it follows that Theorem 7.3 holds whenever N(2m+2)N-(2m+2) is a non-negative integer. This completes the proof.
@@
Acknowledgments. Part of this research was performed while the first author was visiting the Institute for Pure and Applied Mathematics (IPAM), which is supported by the National Science Foundation (Grant No. DMS-1925919). He would like to thank the program organizers and IPAM staff members for their hospitality during his stay in Los Angeles. The second author thanks the Simons Foundation for supporting in part his work by Simons Foundation Collaboration Grant 710477.